# Towards Small Language Models for Security Query Generation in SOC Workflows

### v1

@@ 4.2. RQ2: SLMs with NL2KQL
In this research question, we determine how LLMs and SLMs perform within the NL2KQL system. We reimplement NL2KQL’s architecture with some substitutions to develop a system that can be used to generate KQL queries, and first test this on the four before-mentioned LLMs. NL2KQL selects a certain subset of tables and their respective schema depending on the embedding of the natural language query, and its similarity to queries stored in a few-shot synthetic database, a database that contains realistic, example NLQ-KQL pairs. Furthermore, we test the NL2KQL configuration on the five before-mentioned SLMs. The full prompt used in this system can be referenced within the original NL2KQL paper (nl2kql). These results are shown in Table 2.
From the initial results shown from this table, NL2KQL when partnered with SLMs improves nearly all measured metrics when compared to the initial Zero-Shot approach. Likewise, LLMs show improvement especially in developing semantically correct queries when partnered with NL2KQL. Although SLMs can produce syntactically correct KQL queries in the Zero-Shot configuration, the model struggles to identify the correct tables to use given an NLQ. However, when NL2KQL is utilized to augment the SLM’s knowledge, the SLMs are able to identify proper tables and columns that are needed to not only produce a syntactically correct query but a semantically correct query as well. However each model does incur a greater cost as opposed to utilizing the Zero-Shot approach, and also incurs more costs with respect to input and output tokens. Furthermore, the average latency per query when utilizing the NL2KQL approach is higher compared to the Zero-Shot approach; this is likely due to the higher number of tokens that are given to the model at inference time. We leave the NL2KQL main Avg. Latency and Total Cost for NL2KQL, as these values were not provided in (nl2kql).
@@ 4.3. RQ3: Different Prompt Schemes
In this research question, we seek to understand how different prompting schemes affect NL2KQL with SLMs. In order to revise the prompt to increase the syntactic and semantic score, it is important to understand the syntactic and semantic errors that are generated from using the DeepSeek Coder 6.7B Instruct and the NL2KQL configuration. For this reason, we employ Microsoft’s KQL code parser (Kusto-Query-Language) to first determine the types of syntactic errors that are generated from this scenario. Table 4 in Appendix B lists some of the most common syntactic errors returned from the KQL queries.
@@ 4.4. RQ4: LoRA Fine Tuning
It is possible that due to the nature of reasoning SLMs, supervised fine-tuning alone may not be sufficient in order to train SLMs. When fine-tuning SLMs to improve phishing email detection, researchers found that directly fine-tuning models small reasoning models such as Llama-3.2-3B-Instruct and Qwen-2.5-1.5B-Instruct on phishing emails and respective labels yielded poor results when compared to vanilla prompting strategies (Improving_Phishing_Email_Detection). For this reason, we also attempt a CoT approach, and augment each NLQ–KQL pair with a short explanation generated by the teacher (Gemini 2.0 Flash) of how the KQL result was generated. This provides the SLM model the proper reasoning behind a KQL query result in addition to the actual KQL answer. As outlined in Table 11, since $\alpha=8$, $r=1$, and LoRA dropout = 0.2 minimized validation set loss to 0.0104, we use the fine-tuned model with these parameters. These results are outlined in Table 3 under Advanced Configurations. The CoT Fine-Tuning approach is comparable with the supervised fine-tuning approach in producing syntactically and semantically correct KQL queries. However, it does not surpass the different prompting strategies that have been attempted.
@@ 5. Discussion & Limitations
Generalizability to Languages. The framework extends beyond SQL to EQL(Elastic_Query_Language_Documentation) and SPL(Search_Manual) with straightforward changes. We treat EQL indexes/index patterns and SPL index–sourcetype pairs or CIM datamodels as “tables”; build schemas from Elasticsearch mappings (typed fields, keyword variants, ECS when available) or from Splunk knowledge objects and tstats introspection; then retrieve top-t sources and top-v fields. Few-shot selection is unchanged; examples are serialized in native syntax (EQL event/sequence with where, by, with maxspan; SPL search, where, stats, or tstats when a datamodel applies). Prompts add guardrails: EQL uses canonical timestamp comparisons and restricts fields to chosen indexes; SPL specifies earliest/latest in the search command, prefers tstats for datamodel fields, and avoids table or timechart unless requested. Validators swap accordingly: EQL uses a dry-run parse plus field-existence and legal sequence/maxspan/by checks; SPL uses the parser with per-stage field checks and tstats enforcement. Metrics adapt mechanically: table scores → index or index–sourcetype scores; filter scores = Jaccard over fields/literals in where or stats by, penalizing out-of-schema fields; EQL sequences must have correct stage order. To manage tokens, we cap context to top-t sources and top-v fields and include only relevant operator cheat sheets (e.g., EQL sequence syntax; SPL stats/tstats).
@@ C.1. RQ6: Generalization
After establishing a two-staged architecture and assessing performance on NL2KQL’s evaluation dataset, we also seek to understand how this two-staged architecture performs on other evaluation datasets. In order to test how this system performs on other schemas, we utilize a different dataset from the original evaluation dataset released from NL2KQL. Therefore, we use the Sentinel-Queries Dataset (reprise99) for evaluation, which consists of 83 open-source GitHub KQL queries. We use the same prompting strategy utilized in the Two-Staged Architecture Solution w/Schema Context.

### v2

@@ 4.2. RQ2: SLMs with NL2KQL
In this research question, we determine how LLMs and SLMs perform within the NL2KQL system. We reimplement NL2KQL’s architecture with some substitutions to develop a system that can be used to generate KQL queries, and first test this on the four before-mentioned LLMs. NL2KQL selects a certain subset of tables and their respective schema depending on the embedding of the natural language query, and its similarity to queries stored in a few-shot synthetic database, a database that contains realistic, example NLQ-KQL pairs. Furthermore, we test the NL2KQL configuration on the five before-mentioned SLMs. The full prompt used in this system can be referenced within the original NL2KQL paper (Tang et al., 2025). These results are shown in Table 2.
From the initial results shown from this table, NL2KQL when partnered with SLMs improves nearly all measured metrics when compared to the initial Zero-Shot approach. Likewise, LLMs show improvement especially in developing semantically correct queries when partnered with NL2KQL. Although SLMs can produce syntactically correct KQL queries in the Zero-Shot configuration, the model struggles to identify the correct tables to use given an NLQ. However, when NL2KQL is utilized to augment the SLM’s knowledge, the SLMs are able to identify proper tables and columns that are needed to not only produce a syntactically correct query but a semantically correct query as well. However each model does incur a greater cost as opposed to utilizing the Zero-Shot approach, and also incurs more costs with respect to input and output tokens. Furthermore, the average latency per query when utilizing the NL2KQL approach is higher compared to the Zero-Shot approach; this is likely due to the higher number of tokens that are given to the model at inference time. We leave the NL2KQL main Avg. Latency and Total Cost for NL2KQL, as these values were not provided in (Tang et al., 2025).
@@ 4.3. RQ3: Different Prompt Schemes
In this research question, we seek to understand how different prompting schemes affect NL2KQL with SLMs. In order to revise the prompt to increase the syntactic and semantic score, it is important to understand the syntactic and semantic errors that are generated from using the DeepSeek Coder 6.7B Instruct and the NL2KQL configuration. For this reason, we employ Microsoft’s KQL code parser (17) to first determine the types of syntactic errors that are generated from this scenario. Table 4 in Appendix B lists some of the most common syntactic errors returned from the KQL queries.
@@ 4.4. RQ4: LoRA Fine Tuning
It is possible that due to the nature of reasoning SLMs, supervised fine-tuning alone may not be sufficient in order to train SLMs. When fine-tuning SLMs to improve phishing email detection, researchers found that directly fine-tuning models small reasoning models such as Llama-3.2-3B-Instruct and Qwen-2.5-1.5B-Instruct on phishing emails and respective labels yielded poor results when compared to vanilla prompting strategies (Lin et al., 2025). For this reason, we also attempt a CoT approach, and augment each NLQ–KQL pair with a short explanation generated by the teacher (Gemini 2.0 Flash) of how the KQL result was generated. This provides the SLM model the proper reasoning behind a KQL query result in addition to the actual KQL answer. As outlined in Table 11, since $\alpha=8$, $r=1$, and LoRA dropout = 0.2 minimized validation set loss to 0.0104, we use the fine-tuned model with these parameters. These results are outlined in Table 3 under Advanced Configurations. The CoT Fine-Tuning approach is comparable with the supervised fine-tuning approach in producing syntactically and semantically correct KQL queries. However, it does not surpass the different prompting strategies that have been attempted.
@@ 5. Discussion & Limitations
Generalizability to Languages. The framework extends beyond SQL to EQL(Elastic, 2018) and SPL(Splunk Inc., 2024) with straightforward changes. We treat EQL indexes/index patterns and SPL index–sourcetype pairs or CIM datamodels as “tables”; build schemas from Elasticsearch mappings (typed fields, keyword variants, ECS when available) or from Splunk knowledge objects and tstats introspection; then retrieve top-t sources and top-v fields. Few-shot selection is unchanged; examples are serialized in native syntax (EQL event/sequence with where, by, with maxspan; SPL search, where, stats, or tstats when a datamodel applies). Prompts add guardrails: EQL uses canonical timestamp comparisons and restricts fields to chosen indexes; SPL specifies earliest/latest in the search command, prefers tstats for datamodel fields, and avoids table or timechart unless requested. Validators swap accordingly: EQL uses a dry-run parse plus field-existence and legal sequence/maxspan/by checks; SPL uses the parser with per-stage field checks and tstats enforcement. Metrics adapt mechanically: table scores → index or index–sourcetype scores; filter scores = Jaccard over fields/literals in where or stats by, penalizing out-of-schema fields; EQL sequences must have correct stage order. To manage tokens, we cap context to top-t sources and top-v fields and include only relevant operator cheat sheets (e.g., EQL sequence syntax; SPL stats/tstats).
@@ 6. Conclusion
7. Acknowledgment
This material is based upon work supported by the National Science Foundation (NSF) under Award Nos. 2339483 and 2530655, and by the Commonwealth Cyber Initiative (CCI).
@@ C.1. RQ6: Generalization
After establishing a two-staged architecture and assessing performance on NL2KQL’s evaluation dataset, we also seek to understand how this two-staged architecture performs on other evaluation datasets. In order to test how this system performs on other schemas, we utilize a different dataset from the original evaluation dataset released from NL2KQL. Therefore, we use the Sentinel-Queries Dataset (Zorich, 2021) for evaluation, which consists of 83 open-source GitHub KQL queries. We use the same prompting strategy utilized in the Two-Staged Architecture Solution w/Schema Context.